The Rolling Dough Dessert Company currently has debt which consists of 8 percent coupon bonds (semi-annual coupon payments) which have a maturity of 14 years and are currently priced at $1,154 per bond. There are 12,000 of these bonds outstanding. The firm also has an issue of 1 million preferred shares outstanding with a market price of $14.00 per share. The preferred shares pay an annual dividend of $0.85. RDDC also has 1.5 million shares of common stock outstanding with a price of $19.00 per share. The firm’s beta is 1.3 and the current risk-free rate is 4%. If the expected return on the market is 9% and RDDC is subject to a 40 percent marginal tax rate, then what is the firm's weighted average cost of capital?

WACC please show all steps and formulas
Answer #1
MV of equity=Price of equity*number of shares outstanding
MV of equity=19*1500000
=28500000
MV of Bond=Par value*bonds outstanding*%age of par
MV of Bond=1000*12000*1.154
=13848000
MV of Preferred equity=Price*number of shares outstanding
MV of Preferred equity=14*1000000
=14000000
MV of firm = MV of Equity + MV of Bond+ MV of Preferred equity
=28500000+13848000+14000000
=56348000
Weight of equity = MV of Equity/MV of firm
Weight of equity = 28500000/56348000
W(E)=0.5058
Weight of debt = MV of Bond/MV of firm
Weight of debt = 13848000/56348000
W(D)=0.2458
Weight of preferred equity = MV of preferred equity/MV of firm
Weight of preferred equity = 14000000/56348000
W(PE)=0.2485
Cost of equity
As per CAPM
Cost of equity = risk-free rate + beta * (expected return on the market - risk-free rate)
Cost of equity% = 4 + 1.3 * (9 - 4)
Cost of equity% = 10.5
Cost of debt
                  K = Nx2
Bond Price =∑ [(Semi Annual Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                  K =14x2
1154 =∑ [(8*1000/200)/(1 + YTM/200)^k]     +   1000/(1 + YTM/200)^14x2
                   k=1
YTM = 6.3257157511
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 6.3257157511*(1-0.4)
= 3.79542945066
cost of preferred equity
cost of preferred equity = Preferred dividend/price*100
cost of preferred equity = 0.85/(14)*100
=6.07
WACC=after tax cost of debt*W(D)+cost of equity*W(E)+Cost of preferred equity*W(PE)
WACC=3.8*0.2458+10.5*0.5058+6.07*0.2485
WACC =7.75%